OpenStudy (anonymous):

standard numeral for: (2x10 to the 4th power)+(3x10 to the 2nd power)+(5x10 to the 1st power)+(1x10 to the negative 1st power)

OpenStudy (phi):

add n zeros to the number out front to change to standard decimal example 2*10^2 = 200 (2 with 2 zeros added) for a negative power, add a decimal point and 1 less zero example: 2*10^-2 = 0.01 for your problem, change each to a decimal and add them up

OpenStudy (anonymous):

how does: 20,350.1 sound